OpenStudy (anonymous):

given f(4)=-23 and f(9) = -58, for an arithmetic sequence, find the general term. help please!

OpenStudy (jhannybean):

Arithmetic Progression follows the format \[\large a_{n}=a_{1}+(n-1)d\]

OpenStudy (anonymous):

Very true. Let's get creative. Normally we work with a1, but let's imagine a4 is the first term and a9 is an. What would n be in that case?

OpenStudy (jhannybean):

Oh? Yeah that's somewhat similar to how I solved it...

OpenStudy (anonymous):

Once we know n, we can find d. We can get a1. We can get the general formula.

OpenStudy (anonymous):

How far apart are a4 and a9? That amt takes the place of (n-1).

OpenStudy (anonymous):

so in this case An would be -58?

OpenStudy (jhannybean):

Oh, i see.

OpenStudy (jhannybean):

5-1...ah.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

but why does n become 9?

OpenStudy (jhannybean):

No no, it's not 9.

OpenStudy (jhannybean):

I wish d ended up a pretty number. Hmm....

OpenStudy (anonymous):

We have to look at what the formula really means. It says: an = a term a1 = a term before an (n-1) = how far apart an and a1 are d = how far apart consecutive terms are

OpenStudy (anonymous):

Okay, so that means its-58= -23+(5-1)d

OpenStudy (anonymous):

the question asks to solve for the general term but in this case we find d?

OpenStudy (anonymous):

Not 5-1, but acutally 5. See term 4 and term 9 are 5 places apart.

OpenStudy (jhannybean):

\[\large a_{n}=a_{1}+(n-1)d\] There taking this formula and @mrbarry 's analysis of the arithmetic progression process, we can say that the distance between f(9) and f(4) is 5 numbers. So 5 = n. we're given f(4) = -23 so a1 = -23 we're given up to the number we want to find up to (an) and that's represented by f(9) = -58 we're trying to reach -58.

OpenStudy (anonymous):

@Jhannybean are you comfortable with my generalization of the formula?

OpenStudy (anonymous):

oh gosh, i still dont understand how to solve this question

OpenStudy (jhannybean):

And yes I am @mrbarry

OpenStudy (anonymous):

so substitution and elimination are taking place? -23 = a + d(4-1) -58 = a + d(9-1)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@talha111 remain calm. This is one of those situations where we have to WORK our way to the solution. To get from term 4 to term 9 there are five jumps. To get from -23 to -58 you move -35. Each jump is worth -35/5 = -7 (Now we have d) We can go back to the original equation and use term 4 or term 9 for an. We know all but a1 so now we can find it. a1 =

OpenStudy (anonymous):

Arithmetic sequences are linear functions based on a domain of the natural numbers! You are a clever one.

OpenStudy (jhannybean):

Haha. deleted it because I didn't want to confuse her.

OpenStudy (anonymous):

Are you with us talha? It's annoying when the teachers talk to each other instead of you.

OpenStudy (anonymous):

Im here, its just that ive found a different way to solve the question, using elimination/substitution

OpenStudy (anonymous):

which is what ive been doing for the past couple on minutes

OpenStudy (anonymous):

Good stuff. What did you find? You will need to substitute your a1 and d into a generic equation to get the answer we've sought.

OpenStudy (anonymous):

so far i found d, which is 2.185

OpenStudy (anonymous):

and ive just found a, which is -29.55

OpenStudy (anonymous):

d won't be anything but an integer. Check again.

OpenStudy (anonymous):

a1 is an integer also.

OpenStudy (anonymous):

oh, my bad, ithink ive gotten the general term after fixing it, and thats an=1280(1/2)^n-l

OpenStudy (anonymous):

hopefully this is correct

OpenStudy (jhannybean):

i got an interesting way of looking at the formula. if you didn't understand why (n-1) = 5 and not (9-1) you have n which is the 9th term you want to reach then you have n-1 which includes all the preceding terms. n = 9 n-1 = 4 ((n)- (n-1)) =(n-1) = 9 - 4 = 5 Therefore, you replace (n-1) with 5.